OpenStudy (anonymous):

Write an expansion of the vector x in the given vectors p, q and r. x={-2,0,9}, p={0,-1,2}, q={1,0,-1}, r={-1,2,4}

OpenStudy (anonymous):

Could you be a bit more specific as to what your question is? I'm not sure what it means to expand a vector in vectors.

OpenStudy (anonymous):

It should be expression

OpenStudy (anonymous):

Let a be the number of vector p's you need, let b be the number of vector q's you need, and let r be the number of vector r's you need. Then, x1 = a*(p1) + b*(q1) + c*(r1) x2 = a*(p2) + b*(q2) + c*(r2) x3 = a*(p3) + b*(q3) + c*(r3) solve the system